# LeTO: Learning Constrained Visuomotor Policy with Differentiable Trajectory Optimization

### 30 Jan 2024 v1

@@ Abstract
This paper introduces LeTO, a method for learning constrained visuomotor policy via differentiable trajectory optimization. Our approach uniquely integrates a differentiable optimization layer into the neural network. By formulating the optimization layer as a trajectory optimization problem, we enable the model to end-to-end generate actions in a safe and controlled fashion without extra modules. Our method allows for the introduction of constraints information during the training process, thereby balancing the training objectives of satisfying constraints, smoothing the trajectories, and minimizing errors with demonstrations. This “gray box” method marries the optimization-based safety and interpretability with the powerful representational abilities of neural networks. We quantitatively evaluate LeTO in simulation and on the real robot. In simulation, LeTO achieves a success rate comparable to state-of-the-art imitation learning methods, but the generated trajectories are of less uncertainty, higher quality, and smoother. In real-world experiments, we deployed LeTO to handle constraints-critical tasks. The results show the effectiveness of LeTO comparing with state-of-the-art imitation learning approaches. We release our code at https://github.com/ZhengtongXu/LeTO.
@@ I Introduction
Imitation learning [1] focuses on deriving robot policies from demonstrations. This process can be formulated as a supervised learning task, aiming to learn the mapping between observations and robot actions. Through imitation learning, robots can perform highly complex and diverse tasks.
However, imitation learning has several unique characteristics, such as the presence of multi-modal action distributions and discontinuous functions, sequential correlation, and the demand for precise and smooth trajectories generation. Furthermore, ensuring the safety of robots is very important due to the interaction between robots and the real world. In contrast to optimization-based trajectory generation methods, imitation learning often exhibits greater uncertainty, which can lead to system instability, reduced robustness, and safety concerns. Hence, in comparison to other supervised learning problems, the unique characteristics of imitation learning pose novel and exceptional challenges.
Existing works focus on addressing some of these challenges of imitation learning, such as improving action accuracy by addressing compounding errors [2] and represent multi-modal distributions by implicit policy [3, 4, 5]. However, these methods overlook the critical aspects of safety and smoothness of generated trajectories. For robotic systems, their behaviors need to meet specific constraints to ensure the safety of the system. Moreover, for certain tasks, the robot actions also need to meet specific constraints to ensure their successful completion. In addition, adding explainable constraints for neural network-based algorithms, often called “black boxes”, is much harder than for traditional model-based methods.
@@
Figure 1: In LeTO, we enable the model to end-to-end generate actions in a safe and controlled fashion without extra modules. To the best of our knowledge, LeTO is the first visuomotor imitation learning framework that not only utilizes differentiable optimization but also demonstrates its efficacy in real-world robotic manipulation tasks.
@@
We evaluate LeTO both in simulations (Section IV) and on real robot tasks (Section V) quantitatively, and benchmark with diffusion policy [6], LSTM-GMM [7], and IBC [4]. In the simulated tests, LeTO achieves a success rate comparable to diffusion policy and better than LSTM-GMM and IBC, but the generated trajectories are of less uncertainty, higher quality, and smoother comparing with these methods. On the real robot, we implement constraints-critical manipulation tasks using LeTO. The results highlight LeTO’s strengths compared to the existing imitation learning methods.
@@ II-A Trajectory Optimization
Trajectory optimization is pivotal in generating safe and smooth trajectories for robots and remains an extremely active area of research. CHOMP [8] approaches trajectory optimization by gradient techniques, focusing on optimizing the trade-off between safety and smoothness. TrajOpt [9] uses a sequential convex optimization to generate smooth and collision-free trajecotries and can use naive straight-line initializations that might be in collision. The minimum-snap trajectory optimization [10] uses piecewise polynomials to represent the trajectory and is optimized by quadratic programs. The work in [11] proposes a method for reformulating nondifferentiable collision avoidance constraints into smooth, differentiable constraints and enables real-time optimization-based trajectory planning.
@@ II-C Differentiable Optimization in Robot Learning
For robot navigation and obstacle avoidance, various end-to-end learning frameworks embed with differentiable optimization are proposed, such as using control barrier function [25], gradient-based correction [26], and stacking prediction, planning, and control in a differentiable way [27]. However, all these methods are not suited for performing manipulation tasks. To the best of our knowledge, our work is the first visuomotor imitation learning framework that not only utilizes differentiable optimization but also demonstrates its efficacy in both simulation and real-world robotic manipulation tasks.
@@ III Approach
[…]
Figure 2: Overview of LeTO. We enable the model to end-to-end generate actions in a safe and controlled fashion by integrating a differentiable trajectory optimization layer.
@@ III-A Training Data Sampling
where $\mathbf{y}_{t}=\left[{y}_{t}^{\mathrm{T}},{y}_{t+1}^{\mathrm{T}},\ldots,{y}_{t+T_{p}-1}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{R}^{T_{p}D_{y}}$. $T_{s}$ is the length of each data sampling and $T_{p}$ is the length of the predicted action sequence. The learning objective is minimizing the errors between the predicted $\hat{\mathbf{y}}^{*}_{t}$ and $\mathbf{y}_{t}$ (e.g. using MSE loss to assess the distance between them), as shown in Fig. 2. During policy rollout, the robot will execute the first $T_{a}$-step actions of $\hat{\mathbf{y}}^{*}_{t}$, and then repeat the same process based on new observations. This process is typically called receding horizon planning/control.
@@ III-C Differentiable Trajectory Optimization Layer
By equation (1), the predicted velocity command sequence $\hat{\mathbf{v}}_{t}=\mathbf{S}\hat{\mathbf{y}}_{t}=\left[\hat{v}_{t}^{\mathrm{T}},\hat{v}_{t+1}^{\mathrm{T}},\ldots,\hat{v}_{t+T_{p}-1}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{R}^{T_{p}D_{v}}$, where $\mathbf{S}=\text{blkdiag}\left(S,\dots,S\right)\in\mathbb{R}^{T_{p}D_{v}\times T_{p}D_{y}}$. Then the predicted acceleration sequence and position sequence are
@@ III-D Policy Deployment
$a_{\text{min}}\Delta t$ $\leq\hat{v}_{t}-\hat{v}_{t-1}\leq a_{\text{max}}\Delta t$
where $\hat{v}_{t-1}$ represents the last executed action for the robot at time step $t$ (executed at time step $t-1$). Introducing this new constraint at the test stage will not affect the performance of the trained model, because the sampling of trajectories is random during the training process. Adding this new constraint is transferring the model trained on a discrete set of trajectories into continuous online trajectory optimization.
TABLE II: Trajectory metrics ($10^{-2}$) of the can-mh task. The results correspond to the results in Table I.
@@
TABLE III: Trajectory metrics ($10^{-2}$) of the square-mh task. The results correspond to the results in Table I.
@@
IV Simulation Evaluation
We employed two latest benchmarks used in recent papers for our simulation evaluation: pick and place the can (can task) and grasping and assembling the square nut (square task), as shown in Fig. 3. The results demonstrate that our approach not only achieves success rates comparable to diffusion policy [5], but also significantly surpasses IBC [4] and LSTM+GMM [7]. Furthermore, owing to the presence of constraints and differentiable trajectory optimization, our method exhibits superior performance in terms of generated trajectory quality compared to other methods, which is very important for robot systems.
For training parameters of diffusion policy and LSTM-GMM, we use the default configurations provided in their released code. We also do the same for our realworld evaluation. We will release our code, training data, and checkpoints for reproducing our results.
[…]
We train policies on 3 seeds (42, 142, 242), and test in 25 different initial environments for each seeds (totaling 75 conditions). For each task, we train policies on two different types of datasets: Proficient-Human (ph), and Multi-Human (mh) datasets [7]. We report the success rate (see Table I) of all these experiments. We also report metrics for evaluating the quality of the generated trajectories (see Tables II and III).
@@ IV Simulation Evaluation
We observe that LeTO significantly enhances trajectory smoothness by trajectory optimization while maintaining a consistently high success rate across all simulation tasks, matching the performance of the diffusion policy and surpassing LSTM-GMM and IBC. Interestingly, when applying clipping to constrain the actions generated by the diffusion policy and LSTM-GMM, although this method sometimes increases the smoothness of the trajectory, it results in a reduction in their task success rates.
@@ V Realworld Evaluation
For real robot experiments, we tackled constraints-critical tasks that demand smooth and safe trajectories, which are Move-the-stack (Fig. 4) and Arrange-chopsticks (Fig. 5). We use SpaceMouse to collect human demonstration data. More details of experiments and results can be seen in the attached video.
@@ V-A Move-the-stack
TABLE IV: Metrics for evaluating the Move-the-stack task. “acc-peak” represents the highest acceleration achieved in any of the test trials while “avg-acc-mean/max/std” are the average of mean/max/std values across all test trials.
@@
TABLE V: Metrics for evaluating the Arrange-chopsticks task. “contact-peak” represents the highest contact force achieved in any of the test trials while “avg-contact-max” is the average of max contact forces across all test trials.
@@
TABLE VI: The training time for 20 epochs on the square ph task with a batch size of 64, conducted on a single GeForce RTX 3080. The hyperparameters for diffusion policy and LSTM-GMM are adopted from the default parameters reported in [7, 5], respectively. For the hyperparameters of LeTO, please refer to our open-source code.

### 18 Mar 2024 v2

@@ Abstract
This paper introduces LeTO, a method for learning constrained visuomotor policy via differentiable trajectory optimization. Our approach integrates a differentiable optimization layer into the neural network. By formulating the optimization layer as a trajectory optimization problem, we enable the model to end-to-end generate actions in a safe and constraint-controlled fashion without extra modules. Our method allows for the introduction of constraints information during the training process, thereby balancing the training objectives of satisfying constraints, smoothing the trajectories, and minimizing errors with demonstrations. This “gray box” method marries the optimization-based safety and interpretability with the powerful representational abilities of neural networks. We quantitatively evaluate LeTO in simulation and on the real robot. In simulation, LeTO achieves a success rate comparable to state-of-the-art imitation learning methods, but the generated trajectories are of less uncertainty, higher quality, and smoother. In real-world experiments, we deployed LeTO to handle constraints-critical tasks. The results show the effectiveness of LeTO comparing with state-of-the-art imitation learning approaches. We release our code at https://github.com/ZhengtongXu/LeTO.
@@ I Introduction
Imitation learning [1] focuses on deriving robot policies from demonstrations. This process can be formulated as a supervised learning task, aiming to learn the mapping between observations and robot actions. Through imitation learning, robots can perform highly complex and diverse tasks [2, 3, 4, 5, 6, 7, 8].
However, ensuring the safety of robots is very important due to the interaction between robots and the real world. In contrast to optimization-based trajectory generation methods, imitation learning often exhibits greater uncertainty, which can lead to system instability, reduced robustness, and safety concerns. Hence, in comparison to other supervised learning problems, this pose novel and exceptional challenges in safety for imitation learning.
Existing works focus on different aspects and challenges of imitation learning, such as improving action accuracy by addressing compounding errors [4] and represent multi-modal distributions by implicit policy [9, 10, 2]. However, these methods overlook the critical aspects of safety and smoothness of generated trajectories. For robotic systems, their behaviors need to meet specific constraints to ensure the safety of the system. Moreover, for certain tasks, the robot actions also need to meet specific constraints to ensure their successful completion. In addition, adding explainable constraints for neural network-based algorithms, often called “black boxes”, is much harder than for traditional model-based methods.
@@
Figure 1: In LeTO, we enable the model to end-to-end generate actions in a safe and constraint-controlled fashion without extra modules. To the best of our knowledge, LeTO is the first visuomotor imitation learning framework that not only utilizes differentiable optimization but also demonstrates its efficacy in real-world robotic manipulation tasks.
@@
We evaluate LeTO both in simulations (Section IV) and on real robot tasks (Section V) quantitatively, and benchmark with diffusion policy [11], LSTM-GMM [12], and IBC [10]. In the simulated tests, LeTO achieves a success rate comparable to diffusion policy and better than LSTM-GMM and IBC, but the generated trajectories are of less uncertainty, higher quality, and smoother comparing with these methods. On the real robot, we implement constraints-critical manipulation tasks using LeTO, which highlights LeTO’s strengths.
@@ II-A Trajectory Optimization
Trajectory optimization is pivotal in generating safe and smooth trajectories for robots and remains an active area of research. CHOMP [13] approaches trajectory optimization by gradient techniques, focusing on optimizing the trade-off between safety and smoothness. TrajOpt [14] uses a sequential convex optimization to generate smooth and collision-free trajecotries and can use naive straight-line initializations that might be in collision. The minimum-snap trajectory optimization [15] uses piecewise polynomials to represent the trajectory and is optimized by quadratic programs. The work in [16] proposes a method for reformulating nondifferentiable collision avoidance constraints into smooth, differentiable constraints and enables real-time optimization-based trajectory planning.
@@ II-B Imitation Learning
LeTO focuses on combining explainable and model-based trajectory optimization with imitation learning. In this way, LeTO can be end-to-end trained and generate action in a safe and constraint-controlled fashion without extra modules.
@@ II-C Differentiable Optimization in Robot Learning
The work in [30] proposes a tactile-reactive grasping controller that combines image encoder and differentiable MPC. However, it can not be used for more general policy learning.
For robot navigation and obstacle avoidance, various end-to-end learning frameworks embed with differentiable optimization are proposed, such as using control barrier function [31], gradient-based correction [32], and stacking prediction, planning, and control in a differentiable way [33]. However, all these methods are not suited for performing manipulation tasks.
DiffTOP [34] is a method that integrate differentiable trajectory optimization as the policy representation to generate actions. In integrating imitation learning, DiffTOP focuses on capturing multi-modal distributions. However, it can not generating safe and constrained trajectories. In contrast, LeTO focuses on how the integration of differentiable trajectory optimization into the imitation learning framework enables the generation of trajectories that not only perform tasks effectively but also satisfy constraints.
[…]
Figure 2: Overview of LeTO. We enable the model to end-to-end generate actions in a safe and constraint-controlled fashion by integrating a differentiable trajectory optimization layer.
Another category of methods closely related to LeTO is riemannian motion policy (RMP) [35] and its extensions [36, 37]. Specifically, the works in [36, 37] enhance RMP’s integration into end-to-end robot learning by incorporating automatic differentiation. This enables end-to-end training and inference for policies with RMP structures. Methods based on RMP are particularly focused on generating motion in spaces with high degrees of freedom and nonlinear dynamics. In contrast, LeTO offers a capable form of generating constrained motion at the task space trajectory level with differentiable optimization. This advantage makes it better suited for adaptation to visuomotor policies and demonstrates superior performance in real tasks. To the best of our knowledge, our work is the first visuomotor imitation learning framework that not only utilizes differentiable optimization but also demonstrates its efficacy in both simulation and real-world robotic manipulation tasks.
[…]
Figure 3: Illustration of training data sampling.
@@ III-A Training Data Sampling
where $\mathbf{y}_{t}=\left[{y}_{t}^{\mathrm{T}},{y}_{t+1}^{\mathrm{T}},\ldots,{y}_{t+T_{p}-1}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{R}^{T_{p}D_{y}}$. $T_{s}$ is the length of each data sampling and $T_{p}$ is the length of the predicted action sequence, as shown in Fig. 3. The learning objective is minimizing the errors between the predicted $\hat{\mathbf{y}}^{*}_{t}$ and $\mathbf{y}_{t}$ (e.g. using MSE loss to assess the distance between them), as shown in Fig. 2. During policy rollout, the robot will execute the first $T_{a}$-step actions of $\hat{\mathbf{y}}^{*}_{t}$, and then repeat the same process based on new observations. This process is typically called receding horizon planning/control.
@@ III-C Differentiable Trajectory Optimization Layer
By equation (1), the predicted velocity command sequence
$\hat{\mathbf{v}}_{t}=\mathbf{S}\hat{\mathbf{y}}_{t}=\left[\hat{v}_{t}^{\mathrm{T}},\hat{v}_{t+1}^{\mathrm{T}},\ldots,\hat{v}_{t+T_{p}-1}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{R}^{T_{p}D_{v}},$ (2)
where $\mathbf{S}=\text{blkdiag}\left(S,\dots,S\right)\in\mathbb{R}^{T_{p}D_{v}\times T_{p}D_{y}}$. Then the predicted acceleration sequence and position sequence are
@@
Remark 2 (representational power).
III-D Policy Deployment
[…]
$a_{\text{min}}\Delta t$ $\leq\hat{v}_{t}-\hat{v}_{t-1}\leq a_{\text{max}}\Delta t$
where $\hat{v}_{t-1}$ represents the last executed action for the robot at time step $t$ (executed at time step $t-1$). Adding this new constraint is transferring the model trained on a discrete set of trajectories into continuous online trajectory optimization.
@@ III-D Policy Deployment
IV Simulation Evaluation
We employed two simulation tasks used in [12] for our simulation evaluation: pick and place the can (can task) and grasping and assembling the square nut (square task), as shown in Fig. 4. The results demonstrate that our approach not only achieves success rates comparable to diffusion policy [2], but also significantly surpasses IBC [10] and LSTM+GMM [12]. Furthermore, owing to the presence of constraints and differentiable trajectory optimization, our method exhibits superior performance in terms of generated trajectory quality compared to other methods, which is very important for robot systems.
For training parameters of diffusion policy and LSTM-GMM, we use the default configurations provided in their released code. We also do the same for our realworld evaluation. We have released our code, training data, and checkpoints for reproducing our results.
We train policies on 3 seeds (42, 142, 242), and test in 25 different initial environments for each seeds (totaling 75 conditions). For each task, we train policies on two different types of datasets: Proficient-Human (ph), and Multi-Human (mh) datasets [12]. We report the success rate (see Table I) of all these experiments. We also report metrics for evaluating the quality of the generated trajectories (see Tables II, III, IV, and V).
TABLE II: Trajectory metrics ($10^{-2}$) of the can-ph task. The results correspond to the results in Table I.
DiffusionPolicy [2] $8.23\pm 0.70$ $69.78\pm 2.67$ $8.39\pm 0.58$ $\textbf{1.03}\pm\textbf{ 0.12}$ $12.73\pm 0.78$ $\textbf{1.29}\pm\textbf{0.12}$ 2
LSTM-GMM [12] $9.78\pm 3.54$ $77.49\pm 15.16$ $10.61\pm 3.48$ $\textbf{1.04}\pm\textbf{0.27}$ $24.58\pm 9.73$ $2.71\pm 0.26$ 1
DiffusionPolicy clipping $7.73\pm 0.10$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{3.21}\pm\textbf{ 0.06}$ $3.61\pm 0.47$ $\textbf{9.94}\pm\textbf{0.07}$ $2.45\pm 0.08$ 3
LSTM-GMM clipping $5.61\pm 1.03$ $\textbf{10.00}\pm\textbf{0.00}$ $3.68\pm 0.10$ $1.16\pm 0.13$ $\textbf{9.59}\pm\textbf{0.14}$ $\textbf{1.37}\pm\textbf{0.05}$ 3
LeTO $\textbf{4.36}\pm\textbf{0.10}$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{3.47}\pm\textbf{0.03}$ $1.50\pm 0.21$ $\textbf{9.78}\pm\textbf{0.37}$ $1.73\pm 0.30$ 4
TABLE III: Trajectory metrics ($10^{-2}$) of the can-mh task. The results correspond to the results in Table I.
@@ IV Simulation Evaluation
TABLE IV: Trajectory metrics ($10^{-2}$) of the square-ph task. The results correspond to the results in Table I.
DiffusionPolicy [2] $8.17\pm 0.78$ $70.66\pm 4.49$ $8.65\pm 0.40$ $\textbf{0.95}\pm\textbf{0.11}$ $13.22\pm 0.80$ $\textbf{1.32}\pm\textbf{0.11}$ 2
LSTM-GMM [12] $8.19\pm 2.49$ $68.75\pm 6.52$ $8.76\pm 1.82$ $\textbf{0.91}\pm\textbf{0.08}$ $13.81\pm 3.51$ $\textbf{1.30}\pm\textbf{0.24}$ 2
DiffusionPolicy clipping $6.90\pm 0.75$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{3.36}\pm\textbf{0.06}$ $2.09\pm 0.88$ $\textbf{9.09}\pm\textbf{0.82}$ $1.75\pm 0.59$ 3
LSTM-GMM clipping $\textbf{5.11}\pm\textbf{1.46}$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{3.36}\pm\textbf{0.20}$ $\textbf{0.98}\pm\textbf{0.22}$ $\textbf{9.04}\pm\textbf{0.54}$ $\textbf{1.27}\pm\textbf{0.29}$ 6
LeTO $\textbf{5.08}\pm\textbf{0.33}$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{3.35}\pm\textbf{0.03}$ $1.28\pm 0.17$ $\textbf{9.24}\pm\textbf{0.15}$ $\textbf{1.37}\pm\textbf{0.08}$ 5
TABLE V: Trajectory metrics ($10^{-2}$) of the square-mh task. The results correspond to the results in Table I.
@@
We observe that LeTO significantly enhances trajectory smoothness by trajectory optimization while maintaining a consistently high success rate across all simulation tasks. LeTO achieves a success rate comparable to diffusion policy and surpassing LSTM-GMM and IBC. Interestingly, when applying clipping to constrain the actions generated by the diffusion policy and LSTM-GMM, although this method sometimes increases the smoothness of the trajectory, it results in a reduction in their task success rates.
@@ V Realworld Evaluation
For real robot experiments, we tackled constraints-critical tasks that demand smooth and safe trajectories, which are Move-the-stack (Fig. 5) and Arrange-chopsticks (Fig. 7). We use SpaceMouse teleoperating the end-effector of the robot manipulator to collect human demonstration data. More details of experiments and results can be seen in the attached video.
@@ V-A Move-the-stack
Figure 6: When chopsticks fall, which is a type of OOD data, both the diffusion policy and LeTO can still move the stacked cups and place them on the black board. In contrast, LSTM-GMM often exhibits erroneous behaviors, such as moving to incorrect locations or acting erratically.
TABLE VI: Metrics for evaluating the Move-the-stack task. “acc-peak” represents the highest acceleration achieved in any of the test trials while “avg-acc-mean/max/std” are the average of mean/max/std values across all test trials.
@@
[…]
TABLE VII: Metrics for evaluating the Arrange-chopsticks task. “contact-peak” represents the highest contact force achieved in any of the test trials while “avg-contact-max” is the average of max contact forces across all test trials.
@@
For imitation learning algorithms, the task is challenging because: 1) The policy needs to generate very smooth trajectories. Any roughness in the trajectory can lead to vibrations at the robot’s end-effector, causing the cups to shake. This vibration can accumulate and lead to the chopstick dropping. 2) Since a chopstick is introduced during policy rollout, this out-of-distribution (OOD) observation adds greater uncertainty to the network’s input. This uncertainty may make the network’s output jerkier, resulting in a poorer quality trajectory.
The results are shown in Table VI. From the results, LeTO excels in terms of trajectory smoothness and, correspondingly, also boasts the highest success rate. This demonstrates that due to the presence of end-to-end trajectory optimization, LeTO can robustly maintain its effectiveness even when facing OOD data. Another point worth noting is that when chopsticks fall (which is also a type of OOD data), both the diffusion policy and LeTO can still move the stacked cups and place them on the black board. In contrast, LSTM-GMM often exhibits erroneous behaviors, such as moving to incorrect locations or acting erratically (see Fig. 6 and the attached video). This showcases the robustness of LeTO and the diffusion policy.
V-B Arrange-chopsticks
The challenges of this task include: 1. It is a relative long-horizon task, requiring the policy to have the capability to represent complex tasks. 2. Due to the slender nature of chopsticks, there is a high risk of the robot making forceful contact with the table when attempting to grasp them. For the sake of hardware safety, such collisions should be minimized as much as possible.
The results are shown in Table VII. Both LeTO and diffusion policy have high success rate, but the contact forces in LeTO tests are significantly smaller.
VI Discussion and future work
In this paper, we present LeTO, a framework for learning constrained visuomotor policy with differentiable trajectory optimization. By balancing the objective of minimizing errors with the need to satisfy trajectory constraints, LeTO allows for deployment in tasks where safety and reliability are paramount.
VI-A Hyperparameter, Training Time, and Training Stability
TABLE VIII: The training time for 20 epochs on the square ph task with a batch size of 64, conducted on a single GeForce RTX 3080. The hyperparameters for diffusion policy and LSTM-GMM are adopted from the default parameters reported in [12, 2], respectively. For the hyperparameters of LeTO, please refer to our open-source code.